there were 429 people at a play. admission was $1 for adults and $0.75 for children. the receits were $372.50. how many adults and how many children attended the play?
Answer by CharStar(110)

Adult tickets: x
Adult tickets cost per ticket: $1.00
Total: 1.00x
Children's tickets: (429-x)
Children tickets cost per ticket:$0.75
Total: $0.75(429-x)
1.00x + $0.75(429-x) = 372.50
1.00x + 321.75 - $0.75x = 372.50
0.25x+321.75=372.50
0.25x=$50.75
x=203
Adult tickets = 203.00
Children tickets = $0.75(429-203) = 169.50
Total = 203.00 + 169.50 = 372.50